OpenStudy (anonymous):

at maximum speed, an airplane travels 2100 miles against the wind in 6 hours. flying with the wind, the plane can travel he same distance in 4 hours. what is the speed of the plane with no wind?

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

Recall that d=rt, so 2100=6(v-w)=4(v+w). A little algebra will give us a system of equations v-w=350 v+w=525 Solve this system to get v and w (speed of the jet and speed of the wind).

OpenStudy (anonymous):

I get about 438 for the speed of the airplane, and a wind speed of about 88. Both those numbers are rounded off.